I need some help with this problem
Social Sience
in a town election, the winning candidate had 220 more votes than the loser. if 810 votes were cast in all, how many votes did each candidate receive?

Answer by rapaljer(4551)

Let x = votes cast for the loser
x+ 220 = votes cast for the winner
Total votes cast= 810

x + x+ 220 = 810
2x + 220 = 810
2x+220-220= 810-220
2x= 590
x= 295 votes for the loser
x+ 220 = 515 votes for the winner

Check: 295+ 515 = 810 total votes cast

R^2